# Partial Search in a Frozen Network is Enough to Find a Strong Lottery Ticket

### v2

@@ 4.4 Accuracy to Model Memory Size Trade-Off
Empirically, it is known that SLTs within a SKC initialized dense network achieve better performance than with continuous random weights [30, 26, 21, 38]. Our results show that such a trend can also be observed in frozen source networks. Nonetheless, in all source networks, we find that this improvement is smaller the wider the source network is, suggesting that the requirement for source networks of larger width is weaker in the case of binary weights.
4.5 ImageNet Experiments
Figure 9: Freezing enhances the accuracy-to-model size trade-off even on a practical setting with large-scale dataset and models.
Finally, we evaluate our method using larger models on a large-scale dataset: deeper and wider ResNets on ImageNet. Since SLTs with $80\%$ sparsity achieve the highest accuracy in a dense source ResNet-5050{50}50 (for details, see Figure 10 in Appendix B), we compare the three methods using $80\%$ SLT sparsity. Figures 2 and 9 compare the accuracy-to-search space and accuracy-to-model size trade-offs, respectively, in finding the $80\%$ sparsity SLTs between the proposed and conventional methods with SKC using ImageNet. In both of the figures, such SLTs searched from non-frozen (blue), $70\%$ pruned (orange), and $65\%$ pruned and $5\%$ locked ($70\%$ frozen, green) are compared. Despite the more challenging setting and the significant $70\%$ search space reduction, our method finds SLTs that are substantially more accurate than other methods for the same search space size (Figure 2) and for the same model memory size (Figure 9). This result demonstrates that the effective combination of parameter pruning and locking at initialization can improve the SLT search efficiency even on large-scale datasets and models. Notably, Figure 9 reveals that the proposed method enhances the accuracy-to-model size trade-off, validating our claim it could be deployed on specialized inference hardware, such as the one like Hirose et al. [14], to enhance the efficiency of practical SLT applications.
5 Limitations
Our results have the following limitations: 1) Although the proposed method can effectively reduce the search space for training and the memory model size for inference in principle, actual hardware implementation remains for future work. The previous work of the SLT-specialized neural engine, Hiddenite [14], would be a promising direction for such future implementation. 2) This paper provides theoretical support for the existence of SLTs in frozen networks, but it does not provide their superiority in the approximation capability compared to random pruning, a special case of random freezing, which may depend on the sparsity nature of target networks. Also, the theoretical consideration of the appropriate proportion of these ratios is left for future work. 3) Even though we demonstrate our method on various model architectures including CNNs, ResNet families and GIN, following previous SLT work [41, 30, 21, 17, 38], we have to leave it for future work to apply our method to Transformers [35] because SLT itself has not yet been established for Transformers.
6 Conclusion
This paper capitalizes on the fact that pruning a randomly weighted network at initialization reduces the SLT search space, but identifies that doing so limits the search to a suboptimal sparsity region. This problem is tackled by freezing (i.e., pruning or locking) some parameters at initialization, excluding them from the search. Freezing allows to search for SLTs efficiently in the optimal sparsity range, while further reducing the search space during training and the model size during inference. Experimental results show that SLTs found in frozen networks improve the accuracy-to-search space and accuracy-to-model size trade-offs compared to SLTs found in dense [30, 33] or sparse source networks [10]. Interestingly, although for weight training random parameter locking has only been found useful for reducing accuracy degradation in network compression [41, 36], we identify scenarios in SLT training where they can be used for raising accuracy. Our method can be interpreted as being capable of generating more useful SLT information from a random seed than previous methods, offering an opportunity for reducing off-chip memory access in specialized SLT accelerators [14]. We hope this work will serve as a base for raising the efficiency of SLT training and inference.
Acknowledgments and Disclosure of Funding
This work was supported in part by JSPS KAKENHI Grant Number JP23H05489.
This section describes the proofs of the lemma and theorem introduced in the manuscript. For reference, we present a theorem presented by Gadhikar et al. [10] in advance.
Theorem A.1 (SLT Existence in Sparse Networks [10]).
Let input data $\mathcal{D}$ , a target network $f_{T}$ with depth $L$, and a source network $f_{S}$ with depth $L+1$, parameters $\bm{\theta}_{S}$, and edge probabilities $p_{l}$ be given. Assume that these networks have ReLU activation function, and each element of $\bm{\theta}_{S}$ is uniformly distributed over $[-1,1]$. Then, with probability at least $1-\delta$, there exists a binary mask $\bm{m}_{S}$ so that each output component $i$ is approximated as $\max_{\bm{x}\in\mathcal{D}}\left\|f_{T,i}(\bm{x})-f_{S,i}(\bm{x};\bm{\theta}_{S}\odot\bm{m}_{S})\right\|<\epsilon$ if
$n_{S,l}\geq C\frac{n_{T,l}}{\log(1/1-p_{l+1})}\log\left(\frac{1}{\min\{\epsilon_{l},\delta/\rho\}}\right)$ (2)
for $l\geq 1$,where
$\epsilon_{l}$ $\coloneqq\frac{\epsilon}{n_{T,L}L}\left[(1+B_{l-1}\left(1+\frac{\epsilon}{L}\right)\prod_{k=l+1}^{L-1}\left(\left\|\bm{\theta}_{T}^{(l)}\right\|_{\infty}+\frac{\epsilon}{L}\right)\right]^{-1},$ (3)
$B_{l}$ $\coloneqq\sup_{\bm{x}\in\mathcal{D}}\left\|\bm{x}_{T}^{(l)}\right\|_{1},$ (4)
$\rho$ $=\frac{CN_{T}^{1+\gamma}}{\log\left(1/(1-\min_{l}\{p_{l}\})\right)^{1+\gamma}}\log\left(\frac{1}{\min\{\min_{l}\{\epsilon_{l}\},\delta\}}\right)$ (5)
for any $\gamma\geq 0$. Moreover, we need to satisfy
$n_{S,0}\geq Cd\log\left(\frac{1}{\min\left\{\epsilon_{1},\delta/\rho\right\}}\right),$ (6)
where $C>0$ denotes a generic constant that is independent of $n_{T,l},L,p_{l},\delta$, and $\epsilon$.
We first extend the lemma of the subset-sum approximation shown by Lueker [24] to the case where some random variables are always included in the subset-sum.
Lemma A.2 (Subset-Sum Approximation in Randomly Locked Networks).
Let $X_{1},\cdots,X_{n}\sim U(-1,1)$ be independent, uniformly distributed random variables, and $M_{1},\cdots,M_{n}\sim\mathrm{Ber}(q)$ be independent, Bernoulli distributed random variables with $q\in(0,1)$. Let $\varepsilon,\delta>0$. Then, with probability at least $1-\delta$, for any $z\in[-1,1]$, there exists indices $I\subset\{1,\cdots,n\}$ such that $|z-\sum_{i=1}^{n}M_{i}X_{i}-\sum_{i\in I}(1-M_{i})X_{i}|\leq\varepsilon$ if
$n\geq C\log\left(\frac{8}{\varepsilon\delta}\right),$ (7)
where $C>0$ is a constant.
Proof.
$\mathbb{P}\left(|m-qn|\leq\varepsilon_{M}n\right)\geq 1-2\exp\left(-\frac{\varepsilon_{M}^{2}n}{2q(1-q)}\right)$ (8)
for $\varepsilon_{M}>0$. Thus, if we set $n\geq\frac{2q(1-q)\log(2/\delta)}{\varepsilon_{M}^{2}}$ and $\varepsilon_{M}:=\frac{\min(q,1-q)}{2}$, we have
$(2q-\beta)n\leq m\leq\beta n$ (9)
To approximate $z\in[-1,1]$, we can directly apply Corollary 2.5 from Lueker [24]:
$\mathbb{P}\left(\forall z\in\left[-1,1\right],\exists I\subset\{1,\cdots,n_{1}\}\text{ s.t. }|z-\sum_{i\in I}X^{\prime}_{i}|\geq\varepsilon\right)\geq 1-\delta$ (11)
whenever $n_{1}\geq C\log(2/\varepsilon\delta)$.
From the proof of Corollary 3.1 in Lueker [24], for any $\gamma\in(0,\frac{1}{4})$, we know that
$\mathbb{P}\left(\forall z\in\left[-\gamma n_{2},\gamma n_{2}\right],\exists I\subset\{1,\cdots,n_{2}\}\text{ s.t. }|z-\sum_{i\in I}X^{\prime\prime}_{i}|\geq\varepsilon\right)$ (13)
$\geq 1-\delta-2\exp\left(-\frac{(1-4\gamma)^{2}n_{2}}{64}\right)$ (14)
whenever $n_{2}\geq C\log(\frac{2}{\varepsilon\delta})$. Thus if
$n_{2}\geq\max\left(C\log(\frac{2}{\varepsilon\delta}),\frac{C^{\prime}}{(1-4\gamma)^{2}}\log(\frac{2}{\delta})\right)$ (15)
Now we assume $(\alpha\beta/\gamma)n\leq n_{2}$. For example, if we set $\alpha=\frac{1}{10}$ and $\gamma=\frac{1}{5}$, the assumption is satisfied when we split $\{(1-M_{i})X_{i}\}$ by $n_{1}=n_{2}=\frac{n}{2}$ in (10). Then the approximation (13) can be applied to $z=\sum_{i=1}^{n}M_{i}X_{i}$ with high probability since $|\sum_{i=1}^{n}M_{i}X_{i}|\leq\alpha m\leq\alpha\beta n\leq\gamma n_{2}$ holds.
By combining (11), (12), (15) with the fact that $n=m+n_{1}+n_{2}$, and replacing $\delta$ with $\delta/4$, we obtain the desired results. ∎
By similar arguments as in Lueker [24] or Pensia et al. [28], we can easily generalize Lemma A.2 to the case where the distribution followed by $X_{i}$ contains the uniform distribution. Also, by combining our proof and the proof of Lemma 2.4 in Gadhikar et al. [10], we can prove the following extension of Lemma A.2:
Lemma A.3 (Subset-Sum Approximation in Frozen Networks.).
Let $X_{1},...,X_{n}$ be independent, uniformly distributed random variables so that $X\sim U(-1,1)$, $M_{1},...,M_{n}$ be independent, Bernoulli distributed random variables so that $M_{i}\sim Ber(p)$ for $p\in(0,1)$, and $M^{\prime}_{1},...,M^{\prime}_{n}$ be independent, Bernoulli distributed random variables so that $M^{\prime}_{i}\sim Ber(q)$ for $q\in(0,1)$. Let $\epsilon,\delta\in(0,1)$ given. Then for any $z\in[-1,1]$ there exists a subset $I\subseteq\{1,...,n\}$ so that with probability at least $1-\delta$ we have $\left|z-\sum_{i=1}^{n}M_{i}M^{\prime}_{i}X_{i}-\sum_{i\in I}M_{i}\left(1-M^{\prime}_{i}\right)X_{i}\right|\leq\epsilon$ if
$n\geq C\log\left(\frac{C^{\prime}}{\varepsilon\delta}\right),$ (16)
where $C,C^{\prime}>0$ is a constant.
Theorem A.4 (SLT Existence in Frozen Networks).
Let $\mathcal{D}$ be input data, $f_{T}$ be a target network with depth $L$, and $f_{S}$ be a source network with depth $L+1$, $\bm{\theta}_{S}$ be a parameter of $f_{S}$, and $p_{l},q_{l}\in(0,1)$. Assume that these networks have ReLU activation function, and each element of $\bm{\theta}_{S}$ is uniformly distributed over $[-1,1]$. Also assume that $\bm{\theta}_{S}$ is randomly pruned and locked with pruning ratio $p_{l}$ and locking ratio $q_{l}$ for each $l$-th layer. Then, with probability at least $1-\delta$, there exists a binary mask $\bm{m}_{S}$ so that each output component $i$ is approximated as $\max_{\bm{x}\in\mathcal{D}}\left\|f_{T,i}(\bm{x})-f_{S,i}(\bm{x};\bm{\theta}_{S}\odot\bm{m}_{S})\right\|<\epsilon$ if
$n_{S,l}$ $\geq C_{1}\log\left(\frac{C_{2}}{\min\{\epsilon_{l},\delta/\rho\}}\right)n_{T,l},$ $(l\geq 1)$ (17)
$n_{S,0}$ $\geq C_{3}\log\left(\frac{C_{4}}{\min\left\{\epsilon_{1},\delta/\rho\right\}}\right)d$ $(l=0)$ (18)
where $C_{i}>0$ are constants that include $p_{l},q_{l}$, and $\rho,\epsilon_{l}$ are as defined in Theorem A.1.
Proof.
By replacing Lemma 2.4 in Gadhikar et al. [10] with our Lemma A.3, we obtain the desired result as an extension of Theorem 2.5 in Gadhikar et al. [10]. ∎
Appendix B Additional Experimental Results
B.1 Accuracy of SLT within Dense ResNet-50
This section introduces the preliminary experiment used for determining the main experimental setup. Figure 10 compares the accuracy of SLTs within a dense ResNet-5050{50}50 source at different sparsity using ImageNet. The SLT with $80\%$ sparsity is the most accurate, reaching $66.8\%$ accuracy.
B.2 Result Analysis
Table 1 summarizes the presented results and compares the proposed method to weight training and SLT search in dense and sparse source networks for the same network architecture and SLT sparsity.
In the case of Conv6, compared to weight training our method provides reductions of 2×2$\times$2 × and 66.4×66.4$\times$66.4 × of the search space and model size, respectively, in exchange of a small accuracy drop. Compared to the SLT found in the sparse source network, the SLT found by partial search achieves much higher accuracy.
[…]
Interestingly, the SLT found in the frozen GIN model achieves comparable accuracy to the trained-weight network, even though the model size and search space are reduced by 40.3×40.3$\times$40.3 × and 1.7×1.7$\times$1.7 ×, respectively. The SLT found within a dense network also achieves comparable accuracy to the trained-weight network, but the SLT found by our method is smaller. Generally, graph neural networks (GNNs) suffer from a generalization performance degradation due to the over-smoothing problem [20]. Searching for SLTs within dense GNNs has been shown to mitigate the over-smoothing problem [17] and achieve higher accuracy. As our method is similarly accurate compared to dense networks, it is possible that parameter freezing offers an even stronger solution to the over-smoothing problem.
Although the ImageNet experiment falls into the scenario where a relatively sparse SLT is more accurate (like the ResNet-1818{18}18 experiment on CIFAR-1010{10}10), the SLT found in a sparse network is significantly less accurate. On the other hand, the inclusion of parameter locking shows a significant improvement in SLT accuracy. These results suggest that the number of highly accurate SLT patterns decreases as the difficulty of the problem increases, and the way the strategy of search space reduction affects the performance of the found SLTs more severely. As also shown in Figure 2 in the main manuscript, the SLT found in the frozen Wide ResNet-5050{50}50 achieves $1.4$ points higher accuracy than the SLT found in the dense ResNet-3434{34}34, even though the search space is smaller, indicating superiority of our method in improving the accuracy-to-search space trade-off.
@@ B.2 Result Analysis
Weight Training KU - - - $87.1$ $2.26$ M $8.63$
SLT (Dense) SKC $50$ 00 00 $86.2$ $2.26$ M $0.27$
SLT (Sparse) SKC $50$ $45$ 00 $66.7$ $1.24$ M $0.15$
SLT (Frozen) SKC $50$ $25$ $25$ $84.8$ $1.13$ M $0.13$
@@
Weight Training KU - - - $92.4$ $11.16$ M $42.63$
SLT (Dense) SKC $90$ 00 00 $91.0$ $11.16$ M $1.37$
SLT (Sparse) SKC $90$ $85$ 00 $89.9$ $1.67$ M $0.24$
SLT (Frozen) SKC $90$ $82.5$ $2.5$ $89.8$ $1.67$ M $0.24$
SLT (Frozen) SKC $90$ $85$ $5$ $88.9$ $1.12$ M $0.17$
@@
Weight Training KU - - - $70.1$ $0.38$ M $1.452$
SLT (Dense) SKC $20$ 00 00 $70.0$ $0.38$ M $0.054$
SLT (Sparse) SKC $20$ $15$ 00 $62.3$ $0.32$ M $0.047$
SLT (Frozen) SKC $20$ 00 $40$ $69.2$ $0.23$ M $0.036$
@@
Weight Training KU - - - $74.4$ $25.50$ M $97.49$
SLT (Dense) SKC $80$ 00 00 $66.8$ $25.50$ M $3.24$
SLT (Sparse) SKC $80$ $70$ 00 $47.1$ $7.65$ M $1.11$
SLT (Frozen) SKC $80$ $65$ $5$ $55.2$ $7.65$ M $1.11$
@@
SLT (Dense) SKC $80$ 00 00 $62.5$ $21.78$ M $2.66$
SLT (Sparse) SKC $80$ $70$ 00 $40.7$ $6.53$ M $0.84$
SLT (Frozen) SKC $80$ $65$ $5$ $51.2$ $6.53$ M $0.84$
@@
SLT (Dense) SKC $80$ 00 00 $54.0$ $11.68$ M $1.43$
SLT (Sparse) SKC $80$ $70$ 00 $25.2$ $3.50$ M $0.45$
SLT (Frozen) SKC $80$ $65$ $5$ $39.4$ $3.50$ M $0.45$
@@
SLT (Dense) SKC $80$ 00 00 $70.8$ $68.81$ M $8.46$
SLT (Sparse) SKC $80$ $70$ 00 $57.7$ $20.64$ M $2.72$
SLT (Frozen) SKC $80$ $65$ $5$ $63.9$ $20.64$ M $2.72$
Appendix C Impact Statements

### v3

@@ 4.3 Accuracy-to-Model Memory Size Trade-Off
Empirically, it is known that SLTs within an SKC-initialized dense network achieve better performance than with continuous random weights (Ramanujan et al., 2020; Okoshi et al., 2022; López García-Arias et al., 2023; Yan et al., 2023). Our results show that such a trend can also be observed in frozen source networks. Nonetheless, in all source networks, we find that this improvement is smaller the wider the source network is, suggesting that the requirement for source networks of larger width is weaker in the case of binary weights.
@@
Weight Training KU - - - $87.1$ $8.63$
SLT (Dense) SKC $50$ 00 00 $86.2$ $0.27$
SLT (Sparse) SKC $50$ $45$ 00 $66.7$ $0.15$
SLT (Frozen) SKC $50$ $25$ $25$ $84.8$ $0.13$
@@
Weight Training KU - - - $92.4$ $42.63$
SLT (Dense) SKC $90$ 00 00 $91.0$ $1.37$
SLT (Sparse) SKC $90$ $85$ 00 $89.9$ $0.24$
SLT (Frozen) SKC $90$ $82.5$ $2.5$ $89.8$ $0.24$
SLT (Frozen) SKC $90$ $85$ $5$ $88.9$ $0.17$
@@
Weight Training KU - - - $70.1$ $1.452$
SLT (Dense) SKC $20$ 00 00 $70.0$ $0.054$
SLT (Sparse) SKC $20$ $15$ 00 $62.3$ $0.047$
SLT (Frozen) SKC $20$ 00 $40$ $69.2$ $0.036$
@@
Weight Training KU - - - $74.4$ $97.49$
SLT (Dense) SKC $80$ 00 00 $66.8$ $3.24$
SLT (Sparse) SKC $80$ $70$ 00 $47.1$ $1.11$
SLT (Frozen) SKC $80$ $65$ $5$ $55.2$ $1.11$
@@
SLT (Dense) SKC $80$ 00 00 $62.5$ $2.66$
SLT (Sparse) SKC $80$ $70$ 00 $40.7$ $0.84$
SLT (Frozen) SKC $80$ $65$ $5$ $51.2$ $0.84$
@@
SLT (Dense) SKC $80$ 00 00 $54.0$ $1.43$
SLT (Sparse) SKC $80$ $70$ 00 $25.2$ $0.45$
SLT (Frozen) SKC $80$ $65$ $5$ $39.4$ $0.45$
@@
SLT (Dense) SKC $80$ 00 00 $70.8$ $8.46$
SLT (Sparse) SKC $80$ $70$ 00 $57.7$ $2.72$
SLT (Frozen) SKC $80$ $65$ $5$ $63.9$ $2.72$
4.4 ImageNet Experiments
Finally, we evaluate our method using larger models on a large-scale dataset: deeper and wider ResNets on ImageNet. Since SLTs with $80\%$ sparsity achieve the highest accuracy in a dense source ResNet-5050{50}50 (for details, see Section B.1), we compare the three methods using $80\%$ SLT sparsity.
Figure 2 compares the accuracy-to-model size trade-off in finding the $80\%$ sparsity SLTs between the proposed and conventional methods with SKC using ImageNet. Despite the more challenging setting and the significant $70\%$ memory size reduction, our method (green) finds SLTs that are more accurate than pre-pruning-only methods (orange) for the same model size. This result demonstrates that the effective combination of parameter pruning and locking at initialization can improve the SLT memory efficiency even on large-scale datasets and models.
4.5 Result Analysis
Table 1 summarizes the presented results and compares the proposed method to weight training and SLT search in dense and sparse source networks for the same network architecture and SLT sparsity. As described in Section 4.3, the model size is calculated as the sum of the unfrozen partial supermask (1 bit/parameter) and the batch normalization parameters (32 bits/parameter). Also, as described in Section 4.2, the target SLT sparsity is fixed to that of the best-performing SLT found in a dense source network, and we determine the pre-pruning ratio in sparse network as $(\textrm{SLT sparsity}-5)\%$ for high compression.
In the case of Conv6, compared to weight training our method provides reductions of 66.4×66.4$\times$66.4 × of the model size, in exchange of a small accuracy drop. Compared to the SLT found in the sparse source network, the SLT found in the frozen source network achieves much higher accuracy.
[…]
Interestingly, the SLT found in the frozen GIN model achieves comparable accuracy to the trained-weight network, even though the model size is reduced by 40.3×40.3$\times$40.3 ×. The SLT found within a dense network also achieves accuracy comparable to that of the trained-weight network, but the SLT found by our method is smaller. Generally, graph neural networks (GNNs) suffer from a generalization performance degradation due to the over-smoothing problem Li et al. (2018). Searching for SLTs within dense GNNs has been shown to mitigate the over-smoothing problem Huang et al. (2022) and achieve higher accuracy. As our method is similarly accurate compared to dense networks, it is possible that parameter freezing offers an even stronger solution to the over-smoothing problem.
Although the ImageNet experiment falls into the scenario where a relatively sparse SLT is more accurate (like the ResNet-1818{18}18 experiment on CIFAR-1010{10}10), the SLT found in a sparse network is significantly less accurate. On the other hand, the inclusion of parameter locking shows a significant improvement in SLT accuracy. These results suggest that the number of highly accurate SLT patterns decreases as the difficulty of the problem increases, and the pruning:locking proportion affects the performance of the found SLTs more severely.
5 Limitations
Our results have the following limitations: 1) Although the proposed method can effectively reduce the model memory size for inference in principle, actual hardware implementation remains for future work. The previous work of the SLT-specialized neural engine, Hiddenite Hirose et al. (2022), would be a promising direction for such future implementation. 2) This paper provides theoretical support for the existence of SLTs in frozen networks, but it does not provide their superiority in the approximation capability compared to random pruning, a special case of random freezing, which may depend on the sparsity nature of target networks. Also, the theoretical consideration of the appropriate proportion of these ratios is left for future work. 3) Even though we demonstrate our method on various model architectures including CNNs, ResNet families and GIN in Section 4, following previous SLT work (Zhou et al., 2019; Ramanujan et al., 2020; López García-Arias et al., 2023; Huang et al., 2022; Yan et al., 2023), we have to leave it for future work to apply our method to Transformers (Vaswani et al., 2017) because SLT itself has not yet been established for Transformers.
6 Conclusion
This paper capitalizes on the fact that pre-pruning a randomly weighted network reduces the supermask memory size, but identifies that doing so limits the search to a suboptimal sparsity region. This problem is tackled by freezing (i.e., pruning or locking) some parameters at initialization, excluding them from the search. Freezing allows to the search for SLTs in the optimal sparsity region, while further reducing the model size. Experimental results show that SLTs found in frozen networks improve the accuracy-to-model size trade-off compared to SLTs found in dense (Ramanujan et al., 2020; Sreenivasan et al., 2022) or sparse networks (Gadhikar et al., 2023). Interestingly, although for weight training random parameter locking has only been found useful for reducing accuracy degradation in network compression (Zhou et al., 2019; Wimmer et al., 2020), we identify scenarios in SLT training where they can be used for raising accuracy. Our method can be interpreted as being capable of generating more useful SLT information from a random seed than previous methods, offering an opportunity for reducing off-chip memory access in specialized SLT accelerators (Hirose et al., 2022). Additionally, the reduced number of parameters to be optimized may be exploited for training cost reduction. Appropriate freezing may improve the SLT search efficiency, including preventing convergence to local low-precision solutions as experimentally confirmed in Section 4.2, and improving convergence speed through reduced parameters to be optimized. Investigations about the SLT search efficiency remain as future work.
Acknowledgments
This work was supported in part by JSPS KAKENHI Grant Number JP23H05489 and JST-ALCA-Next Japan Grant # JPMJAN24F3.
Impact Statements